4v^2+16v+23=0
Complete the square

Answer by ewatrrr(24785)


Hi
Complete the square
4v^2+16v+23=0
4(v^2+4v)+23=0
4[(v+2)^2 - 4]+23=0
4(v+2)^2 - 16 + 23 = 0
4(v+2)^2 + 7 = 0